How are the E and H fields related in a plane electromagnetic wave?
If the amplitude of the electric part So of a plane wave in vacuum is 30volts/m,what will be the amplitude of the associated magnetic part Ho. Calculate the average energy flow of the electromagnetic wave in watt/m^2(7.96×10^-2T,1.2w/m^2)
1
Expert's answer
2021-07-15T13:52:16-0400

An electromagnetic wave consists of both type of field, which are oscillating back and forth. The relationship between electric and magnetic fields is what allows the formation of electromagnetic waves, including light and heat. This relationship is fundamental to the working of the universe in its present form.


"H=\frac{E}{c\mu_0}\\H=\frac{30}{3\cdot10^{8}\cdot 4\pi\cdot10^{-7}}=0.0796\ T\\
S=(30)(0.0796)=1.2\frac{W}{m^2}"